OpenStudy (jb1625):

-1/7 - 2/5 A.-9/35 B.9/35 C.19/35 D.-19/35

OpenStudy (jb1625):

Dude can you please type faster or something?

OpenStudy (anonymous):

So first, setup the equation like so.. \[-\frac{ 1 }{ 7 } - \frac{ 2 }{ 5 }\] Put the fractions over a common denominator of 35. \[-\frac{ 1 }{ 7 }-\frac{ 2 }{ 5 } = \frac{ -5 }{ 35 } + \frac{ 7(-2) }{ 35 }\] Now multiply.. \[\frac{ -5 }{ 35 }+\frac{ -14 }{ 35 }\] Now go ahead and add for your final answer which is.. \[\frac{ -19 }{ 35 }\]

OpenStudy (jb1625):

Ok thank you

OpenStudy (jb1625):

And can you tell me how you wrote in that fraction?

OpenStudy (anonymous):

Wait, my bad.. Not -19/35 it should look more so like this.. and you're welcome :). \[-\frac{ 19 }{ 35 }\]

OpenStudy (anonymous):

Click the "Equation" button below the box you type in. It'll say "Equation, Draw, Attach File", and "Post".

OpenStudy (jb1625):

Oh ok thank you